# Towards an Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning

### v1

@@ Abstract
As a marriage between offline RL and meta-RL, the advent of offline meta-reinforcement learning (OMRL) has shown great promise in enabling RL agents to multi-task and quickly adapt while acquiring knowledge safely. Among which, Context-based OMRL (COMRL) as a popular paradigm, aims to learn a universal policy conditioned on effective task representations. In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified information theoretic framework. Most importantly, we show that the pre-existing COMRL algorithms are essentially optimizing the same mutual information objective between the task variable $\bm{M}$ and its latent representation $\bm{Z}$ by implementing various approximate bounds. Based on the theoretical insight and the information bottleneck principle, we arrive at a novel algorithm dubbed UNICORN, which exhibits remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures, attaining the new state-of-the-art. We believe that our framework could open up avenues for new optimality bounds and COMRL algorithms.
@@ 1 Introduction
The ability to swiftly learn and generalize to new tasks is a hallmark of human intelligence. In pursuit of this high level of artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments. In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data. This meta-RL mechanism has been adopted in many applications such as games (Zhao et al., 2022a; Beck et al., 2023), robotics (Yu et al., 2020a; Kumar et al., 2021) and drug discovery (Feng et al., 2022).
@@
However, for data collection, classical meta-RL usually requires enormous online explorations of the environments (Finn et al., 2017; Rakelly et al., 2019), which is impractical in many safety-critical scenarios such as healthcare (Gottesman et al., 2019) and autonomous driving (Shalev-Shwartz et al., 2016). As a remedy, offline RL (Levine et al., 2020) enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions. Recently, offline meta-RL (OMRL) (Li et al., 2021b; Mitchell et al., 2021; Dorfman et al., 2021) has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure. In the context of supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning (Raffel et al., 2020), meta-training (Finn et al., 2017) and fine-tuning (Hu et al., 2021; Li and Liang, 2021; Lester et al., 2021) of pre-trained large models. We envision it to be a cornerstone of RL foundation models (Bommasani et al., 2021; Reed et al., 2022) in the future.
@@
Along the line of OMRL research, context-based offline meta-reinforcement learning (COMRL) is a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Intuitively, the crux of COMRL is learning effective task representations, and therefore enabling the agent to react optimally and adaptively in various contexts. To achieve this goal, one of the earliest COMRL algorithms FOCAL (Li et al., 2021b) proposes to capture the structure of task representations by distance metric learning. From a geometric perspective, it essentially performs clustering by pushing away latent embeddings of different tasks while pulling together those from the same task, hence ensuring consistent and distinguishable task representations.
@@
Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts (Li et al., 2021a), i.e., when testing on out-of-distribution (OOD) data. Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance (Li et al., 2021a; Gao et al., 2023). To alleviate the problem, follow-up works such as CORRO (Yuan and Lu, 2022) reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$. It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE (Oord et al., 2018) contrastive loss, where the positive and negative pairs are conditioned on identical state-action tuples $(\bm{s},\bm{a})$. Inspired by CORRO, a recently proposed method CSRO (Gao et al., 2023) introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$. By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo (Todorov et al., 2012) benchmarks.
In this paper, following the recent development and storyline of COMRL, we present a novel and unified view of several signature methods including FOCAL, CORRO and CSRO under a information theoretic framework. We first show that these algorithms are essentially optimizing different approximate bounds of the same mutual information objective $I(\bm{Z},\bm{M})$. Then based on our proposed Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning (UNICORN), we arrive at a new COMRL optimization scheme by deriving a unified and generalized task representation learning objective in light of the information bottleneck principle  (Tishby et al., 1999; Tishby and Zaslavsky, 2015). We demonstrate experimentally that the proposed method achieves state-of-the-art in-distribution and OOD generalization performance on a wide range of RL domains, OOD settings, data qualities and model architectures. Our framework potentially provides a principled roadmap to new COMRL algorithms by seeking better optimality bounds/approximations of the objective, as well as new implementations to further combat context shift.
2 Related Works
Offline Meta-Reinforcement Learning.   OMRL solves the problems of generalization and expensive data collection by meta-learning on pre-collected offline datasets. To overcome the overestimation of out-of-distribution actions (Kumar et al., 2019; Levine et al., 2020; Gao et al., 2022), offline RL usually leverages regularizations such as KL-divergence (Jaques et al., 2019), maximum mean discrepancy (MMD) (Kumar et al., 2019) and Wasserstein Distance (Gulrajani et al., 2017) to constrain the distribution mismatch between the behavior and learner policies. Moreover, a general mechanism called behavior regularized actor critic (BRAC) (Wu et al., 2019) involving value penalty and behavior regularization has proven successful in several previous studies (Li et al., 2021b; Gao et al., 2023), and is adopted in our method.
From a meta-learning perspective, the realization of OMRL can be categorized as gradient-based and context-based paradigms. Gradient-based meta-RL methods (Finn et al., 2017; Mitchell et al., 2021) embed an optimization procedure such as gradient descent into the meta-level iteration loops, which are model-agnostic but tend to reach lower asymptotic performance with less computational efficiency in practice. COMRL conditions the models on a contextual representation of MDP (Hallak et al., 2015) generated by recurrent (Wang et al., 2016; Fakoor et al., 2020), recursive (Mishra et al., 2017; Xu et al., 2022), probabilistic (Rakelly et al., 2019; Zintgraf et al., 2019) or deterministic (Li et al., 2021b, a; Yuan and Lu, 2022) task inference modules. At test time, it leverages neural networks as universal value/policy function approximators (Schaul et al., 2015) to perform meta-adaptation by transfer learning of the task representations. Depending on the source of the testing data, fully offline (Li et al., 2021b, a; Yuan and Lu, 2022) and offline-to-online (Zhao et al., 2022b; Pong et al., 2022) COMRL have been investigated. Notwithstanding, both paradigms may suffer significant OOD as the data-collecting policies might drastically differ from the training sets. In this work, we restrict our attention to fully offline COMRL, and demonstrate the remarkable robustness of our proposed framework to context shift in this setting.
@@ 3 Method
Figure 1: Context shift of COMRL in Ant-Dir. Left: Given a task specified by a goal direction (dashed line), the RL agent is trained on data generated by a variety of behavior policies trained on the same task (red). At test time, however, the context might be collected by behavior policies trained on different tasks (blue), causing a context shift of OOD behavior policies (Section 4.4). Middle: Under OOD testing context, UNICORN (red) is more robust than baselines such as FOCAL (green) in terms of navigating the Ant robot along the right direction. Right: Besides behavior policy, the task distribution (e.g., goal positions in Ant) can induce significant context shift (Section 5.2), which is also a challenging scenario for COMRL models to generalize.
3.1 Preliminaries and Problem Statement
We consider MDP modeled as $M=(\mathcal{S},\mathcal{A},T,R,\mathcal{\rho}_{0},\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $T(\bm{s}^{\prime}|\bm{s},\bm{a})$, bounded reward function $R(\bm{s},\bm{a})$, initial state distribution $\rho_{0}(\bm{s})$ and discount factor $\gamma\in(0,1)$. The goal is to find a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ to maximize the expected return. Starting from the initial state, for each time step, the agent performs an action sampled from $\pi$, then the environment updates the state with $T$ and returns a reward with $R$. We denote the marginal state distribution at time $t$ as $\mu_{\pi}^{t}(\bm{s})$. The V-function and Q-function are given by
$V_{\pi}(\bm{s})$ $=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\bm{s}_{t}\sim\mu_{\pi}^{t}(\bm{s}),\bm{a}_{t}\sim\pi}[R(\bm{s}_{t},\bm{a}_{t})]$ (1)
@@ 3.1 Preliminaries and Problem Statement
In this paper, we restrict to scenarios where tasks share the same state and action space while differing in reward and transition functions, with each OMRL task defined as an instance of MDP: $M^{i}=(\mathcal{S},\mathcal{A},T^{i},R^{i},\rho_{0},\gamma)\in\mathcal{M}$, where $\mathcal{M}$ is the set of all possible MDPs to be considered. For each task index $i$, the offline dataset $X^{i}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}$ is collected by a behavior policy $\pi^{i}_{\beta}$. For meta-learning, given a trajectory segment $\bm{c}^{i}_{1:n}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}_{j=1}^{n}$ of length $n$ as the context of $M^{i}$, COMRL employs a context encoder $q_{\bm{\phi}}(\bm{z}|\bm{c}_{1:n}^{i})$ to obtain the latent representation $\bm{z}^{i}$ of task $M^{i}$. If $\bm{z}$ contains enough information about the task identity, COMRL can be treated as a special case of RL on partially-observed MDP (Kaelbling et al., 1998), where $\bm{z}$ is interpreted as a faithful representation of the unobserved state. By conditioning on $\bm{z}$, the learning of universal policy $\pi_{\bm{\theta}}(\cdot|\bm{s},\bm{z})$ and value function $V_{\pi}(\bm{s},\bm{z})$ (Schaul et al., 2015) become regular RL, and optimality can be attained by Bellman updates with theoretical guarantees. To this end, FOCAL (Li et al., 2021b) proposes to decouple COMRL into the upstream task representation learning and downstream offline RL. For the former, FOCAL effectively clusters task embeddings by metric learning:
$\mathcal{L}_{\text{FOCAL}}=$ $\,\,\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{i,j}\text{\LARGE\{}\textbf{1}\{i=j\}||\bm{z}^{i}-\bm{z}^{j}||_{2}^{2}+\textbf{1}\{i\neq j\}\frac{\beta}{||\bm{z}^{i}-\bm{z}^{j}||_{2}^{n}+\epsilon}\text{\LARGE\}}.$ (3)
However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift (Li et al., 2021a). This problem has been formulated in several earliest studies of OMRL.  Li et al. (2020) observed that when there are large divergence between the state-action distributions of the offline datasets, due to shortcut learning (Geirhos et al., 2020), the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.  Dorfman et al. (2021) concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL. An example is illustrated in Figure 1. The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
To mitigate context shift, a subsequent algorithm CORRO (Yuan and Lu, 2022) proposes to optimize a lower bound of $I(\bm{Z};\bm{M})$ in the form of an InfoNCE (Oord et al., 2018) contrastive loss
@@
where $\bm{x}=(\bm{s},\bm{a},\bm{r},\bm{s}^{\prime})$, $\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})$, $h(\bm{x},\bm{z})=\frac{P(\bm{z}|\bm{x})}{P(\bm{z})}\approx\frac{q_{\bm{\phi}}(\bm{z}|\bm{x})}{P(\bm{z})}$, and $\bm{x}^{*}=(\bm{s},\bm{a},r^{*},\bm{s}^{\prime*})$ as a transition tuple generated for task $M^{*}\in\mathcal{M}$ conditioned on the same $(\bm{s},\bm{a})$. To further combat the spurious correlation between the task representation and behavior-policy-induced state-actions, a recently proposed method CSRO (Gao et al., 2023) introduces a CLUB upper bound (Cheng et al., 2020) of the mutual information between $\bm{z}$ and $(\bm{s},\bm{a})$, to regularize the FOCAL objective:
$\mathcal{L}_{\text{CSRO}}$ $=\underset{\bm{\phi}}{\min}\left\{\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}\right\}$ (5)
@@
In the next section, we will show how these algorithms are inherently connected and can be further improved under our proposed unified information theoretic framework.

### v2

@@ Abstract
As a marriage between offline RL and meta-RL, the advent of offline meta-reinforcement learning (OMRL) has shown great promise in enabling RL agents to multi-task and quickly adapt while acquiring knowledge safely. Among which, context-based OMRL (COMRL) as a popular paradigm, aims to learn a universal policy conditioned on effective task representations. In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified framework. Most importantly, we show that the pre-existing COMRL algorithms are essentially optimizing the same mutual information objective between the task variable $\bm{M}$ and its latent representation $\bm{Z}$ by implementing various approximate bounds. Such theoretical insight offers ample design freedom for novel algorithms. As demonstrations, we propose a supervised and a self-supervised implementation of $I(\bm{Z};\bm{M})$, and empirically show that the corresponding optimization algorithms exhibit remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures. This work lays the information theoretic foundation for COMRL methods, leading to a better understanding of task representation learning in the context of reinforcement learning  [^1] The open-source code is at https://github.com/betray12138/UNICORN.git..
@@ 1 Introduction
The ability to swiftly learn and generalize to new tasks is a hallmark of human intelligence. In pursuit of this high level of artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments. In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data. This meta-RL mechanism has been adopted in many applications such as games [1, 2], robotics [3, 4] and drug discovery [5].
@@
However, for data collection, classical meta-RL usually requires enormous online explorations of the environments [6, 7], which is impractical in many safety-critical scenarios such as healthcare [8] and robotic manipulation [9]. As a remedy, offline RL [10] enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions. Recently, offline meta-RL (OMRL) [11, 12, 13] has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure. In the context of classical supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning [14], meta-training [6] and fine-tuning [15, 16, 17] of pre-trained large models. We envision it as a cornerstone of RL foundation models  [18, 19, 20] in the future.
@@
Along the line of OMRL research, context-based offline meta-reinforcement learning (COMRL) is a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Intuitively, the crux of COMRL lies in learning effective task representations, hence enabling the agent to react optimally and adaptively in various contexts. To this end, one of the earliest COMRL algorithms FOCAL [11] proposes to capture the structure of task representations by distance metric learning. From a geometric perspective, it essentially performs clustering by repelling latent embeddings of different tasks while pulling together those from the same task, therefore ensuring consistent and distinguishable task representations.
@@
Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts [21], i.e., when testing on out-of-distribution (OOD) data (Fig. 1). Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance  [21, 22]. To alleviate the problem, follow-up works such as CORRO [23] reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$. It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE [24] contrastive loss, where the positive and negative pairs are conditioned on the state-action tuples $(\bm{s},\bm{a})$. Inspired by CORRO, a recently proposed method CSRO [22] introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$. By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo [25] benchmarks.
Contributions  In this paper, following the recent development and storyline of COMRL, we present a Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning (UNICORN) encompassing pre-existing methods. We first prove that the objectives of FOCAL, CORRO and CSRO operate as the upper bound, lower bound of $I(\bm{Z};\bm{M})$ and their linear interpolation respectively, which provides a nontrivial theoretical unification of these methods.
Second, by the aforementioned insight and an analysis of the COMRL causal structures, we shed light on how CORRO and CSRO improve context-shift robustness compared to their predecessors by trading off causal and spurious correlations between $\bm{Z}$ and input data $\bm{X}$.
Lastly, by examining eight related meta-RL methods (Table 1) concerning their objectives and implementations, we highlight the potential design choices of novel algorithms offered by our framework. As examples, we investigate two instantiated algorithms, one supervised and the other self-supervised, and demonstrate experimentally that they achieve competitive in-distribution and exceptional OOD generalization performance on a wide range of RL domains, OOD settings, data qualities and model architectures. Our framework provides a principled roadmap to novel COMRL algorithms by seeking better approximations/regularizations of $I(\bm{Z};\bm{M})$, as well as new implementations to further combat context shift.
@@ 2 Method
Figure 1: Context shift of COMRL in Ant-Dir. Left: Given a task $M^{i}$ specified by a goal direction (dashed line), the RL agent is trained on data generated by a variety of behavior policies trained on the same task $M^{i}$ (red). At test time, however, the context might be collected by behavior policies trained on different tasks $\{M^{j}\}$ (blue), causing a context shift of OOD behavior policies (Section 3.3). Middle: Against OOD context, UNICORN (red) is more robust than baselines such as FOCAL (green) in terms of navigating the Ant robot towards the right direction. Right: Besides behavior policy, the task distribution (e.g., goal positions in Ant) can induce significant context shift (Section 4.2), which is also a challenging scenario for COMRL models to generalize.
2.1 Preliminaries, Problem Statement and Related Work
We consider MDP modeled as $M=(\mathcal{S},\mathcal{A},T,R,\mathcal{\rho}_{0},\gamma,H)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $T(\bm{s}^{\prime}|\bm{s},\bm{a})$, bounded reward function $R(\bm{s},\bm{a})$, initial state distribution $\rho_{0}(\bm{s})$, discount factor $\gamma\in(0,1)$ and $H$ the horizon. The goal is to find a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ to maximize the expected return. Starting from the initial state, for each time step, the agent performs an action sampled from $\pi$, then the environment updates the state with $T$ and returns a reward with $R$. We denote the marginal state distribution at time $t$ as $\mu_{\pi}^{t}(\bm{s})$. The V-function and Q-function are given by
$V_{\pi}(\bm{s})$ $=\sum_{t=0}^{H-1}\gamma^{t}\mathbb{E}_{\bm{s}_{t}\sim\mu_{\pi}^{t}(\bm{s}),\bm{a}_{t}\sim\pi}[R(\bm{s}_{t},\bm{a}_{t})],$ (1)
@@ 2.1 Preliminaries, Problem Statement and Related Work
In this paper, we restrict to scenarios where tasks share the same state and action space and can only be distinguished via reward and transition functions. Each OMRL task is defined as an instance of MDP: $M^{i}=(\mathcal{S},\mathcal{A},T^{i},R^{i},\rho_{0},\gamma,H)\in\mathcal{M}$, where $\mathcal{M}$ is the set of all possible MDPs to be considered. For each task index $i$, the offline dataset $X^{i}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}$ is collected by a behavior policy $\pi^{i}_{\beta}$. For meta-learning, given a trajectory segment $\bm{c}^{i}_{1:n}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}_{j=1}^{n}$ of length $n$ as the context of $M^{i}$, COMRL employs a context encoder $q_{\bm{\phi}}(\bm{z}|\bm{c}_{1:n}^{i})$ to obtain the latent representation $\bm{z}^{i}$ of task $M^{i}$. If $\bm{z}$ contains sufficient information about the task identity, COMRL can be treated as a special case of RL on partially-observed MDP [26], where $\bm{z}$ is interpreted as a faithful representation of the unobserved state. By conditioning on $\bm{z}$, the learning of universal policy $\pi_{\bm{\theta}}(\cdot|\bm{s},\bm{z})$ and value function $V_{\pi}(\bm{s},\bm{z})$ [27] become regular RL, and optimality can be attained by Bellman updates with theoretical guarantees. To this end, FOCAL [11] proposes to decouple COMRL into the upstream task representation learning and downstream offline RL. For the former, which can be treated independently, FOCAL employs metric learning to achieve effective clustering of task embeddings:
$\mathcal{L}_{\text{FOCAL}}=$ $\,\,\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{i,j}\left\{\mathbbm{1}\{i=j\}||\bm{z}^{i}-\bm{z}^{j}||_{2}^{2}+\mathbbm{1}\{i\neq j\}\frac{\beta}{||\bm{z}^{i}-\bm{z}^{j}||_{2}^{n}+\epsilon}\right\}.$ (3)
However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift  [21]. This problem has been formulated in several earliest studies of OMRL.  Li et al. [28] observed that when there are large divergence of the state-action distributions among the offline datasets, due to shortcut learning [29], the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.  Dorfman et al. [13] concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL. An example is illustrated in Figure 1. The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
To mitigate context shift, a subsequent algorithm CORRO [23] proposes to optimize a lower bound of $I(\bm{Z};\bm{M})$ in the form of an InfoNCE [24] contrastive loss
@@
where $\bm{x}=(\bm{s},\bm{a},r,\bm{s}^{\prime})$, $\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})$, $h(\bm{x},\bm{z})=\frac{P(\bm{z}|\bm{x})}{P(\bm{z})}\approx\frac{q_{\bm{\phi}}(\bm{z}|\bm{x})}{P(\bm{z})}$, and $\bm{x}^{*}=(\bm{s},\bm{a},r^{*},\bm{s}^{\prime*})$ as a transition tuple generated for task $M^{*}\in\mathcal{M}$ conditioned on the same $(\bm{s},\bm{a})$. To further combat the spurious correlation between the task representation and behavior-policy-induced state-actions, a recently proposed method CSRO [22] introduces a CLUB upper bound [30] of the mutual information between $\bm{z}$ and $(\bm{s},\bm{a})$, to regularize the FOCAL objective:
$\mathcal{L}_{\text{CSRO}}$ $=\underset{\bm{\phi}}{\min}\left\{\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}\right\},$ (5)
@@
In the next section, we will show how these algorithms are inherently connected and how their context-shift robustness gets improved incrementally.